# Connectivity Structure and Dynamics of Nonlinear Recurrent Neural Networks

### v1

@@ III.4 Adding sources
Consider a general intensive action $\mathcal{S}[\bm{C};\mathcal{J}]$ where $\bm{C}$ is a collection of fields (e.g., $\bm{C}=\{C^{\phi},\hat{C}^{\phi}\}$), and $\mathcal{J}$ is a source. In Appendix D, we show that
$\left.\frac{\delta^{2}}{\delta\mathcal{J}^{2}}\log Z\right|_{\mathcal{J}=0}\\ =\left[-\frac{\delta^{2}S}{\delta\mathcal{J}^{2}}+\left(\frac{\delta^{2}S}{\delta\mathcal{J}\delta\bm{C}}\right)^{T}\bm{H}^{-1}\frac{\delta^{2}S}{\delta\mathcal{J}\delta\bm{C}}\right]_{\mathcal{J}=0},$ (54)
@@
III.5 Path-integral analysis of the random-mode model
Having demonstrated the path-integral approach to computing two- and four-point functions in well known connectivity models, we now apply this formalism to the random-mode model. In terms of the mode matrices $\bm{L}$ and $\bm{R}$, the path integral is
$Z[\bm{L},\bm{R}]=\int\mathcal{D}\bm{x}\int\mathcal{D}\hat{\bm{x}}\\ \times\exp\left(i\sum_{i}\hat{x}_{i}T[x_{i}]-i\sum_{i,j}\left[\bm{L}\bm{D}\bm{R}\right]_{ij}\hat{x}_{i}\phi_{j}\right).$ (55)
@@ III.5 Path-integral analysis of the random-mode model
$Z[\bm{L},\bm{R}]=\int\mathcal{D}\bm{x}\int\mathcal{D}\hat{\bm{x}}\int\mathcal{D}\bm{z}\int\mathcal{D}\hat{\bm{z}}\\ \times\exp\Big{(}i\sum_{i}\hat{x}_{i}T[x_{i}]+i\sum_{a}z_{a}\hat{z}_{a}\\ -i\sum_{i,a}L_{ia}\hat{x}_{i}z_{a}-i\sum_{i,a}D_{a}R_{ia}\phi_{i}\hat{z}_{a}\Big{)}.$ (57)
@@
$Z$ $=\int\mathcal{D}C^{\phi}\int\mathcal{D}\hat{C}^{\phi}\int\mathcal{D}Q\int\mathcal{D}\hat{Q}\exp\left(-N\mathcal{S}[C^{\phi},\hat{C}^{\phi},Q,\hat{Q}]\right),$ (59)
@@
$W^{x}[Q,\hat{Q}]$ $=\int\mathcal{D}x\int\mathcal{D}\hat{x}\exp\left(i\hat{x}T[x]-\frac{1}{2}\phi\hat{Q}\phi-\frac{1}{2}\hat{x}Q\hat{x}\right),$ (61)
$W^{z}_{D}[C^{\phi},\hat{C}^{\phi}]$ $=\int\mathcal{D}z\int\mathcal{D}\hat{z}\exp\left(i\hat{z}z-\frac{1}{2}z\hat{C}^{\phi}z-\frac{1}{2}D^{2}\hat{z}C^{\phi}\hat{z}\right).$ (62)
@@
III.5.1 Two-point functions
Following the steps outlined at the end of Sec. III.3, we first compute the DMFT by calculating the saddle point. The derivatives of the action are
@@ III.5.1 Two-point functions
Solving for $Q_{\star}(t_{1},t_{2})$ by combining Eq. 69, Eq. 71, and Eq. 73, we obtain $Q_{\star}(t_{1},t_{2})=\alpha r_{2}C^{\phi}_{\star}(t_{1},t_{2}),$ with $r_{n}$ defined by Eq. 25. Consolidating these results, the self-consistent problem that determines $C^{\phi}_{\star}(t_{1},t_{2})$ can be expressed purely in terms of this correlation function as
$T[x](t)$ $=\eta^{x}(t),$
$\eta^{x}$ $\sim\mathcal{GP}(0,\alpha r_{2}C^{\phi}_{\star}),$
$C_{\star}^{\phi}(t_{1},t_{2})$ $=\left\langle\phi(t_{1})\phi(t_{2})\right\rangle_{\star}.$
Comparing to Eqs. 1719, we recover the self-consistent equations of the i.i.d. model with coupling strength
$g_{\text{eff}}=\sqrt{\alpha r_{2}}.$ (74)
Thus, the low-rank coupling structure remains undetectable at the single-neuron level, where neurons experience effectively i.i.d. connections with variance $g_{\text{eff}}^{2}/N$. The structure of the random-mode model emerges only in collective neural activity, as we now demonstrate by calculating the four-point function.
III.5.2 Four-point functions
@@
$\bm{H}=\left(\begin{array}[]{cc|cc}0&\alpha r_{2}&0&-1\\ \alpha r_{2}&-\alpha r_{4}C_{12}^{\phi}&-1&0\\ \hline\cr 0&-1&0&(S_{12}^{\phi})^{*}\\ -1&0&S_{12}^{\phi}&-C_{12}^{\phi}\end{array}\right).$ (75)
@@
The upper-left block, corresponding to the latent variables, is related to the Hessian of an i.i.d. network with linear dynamics and a heterogeneous distribution of single-neuron scale factors $D_{a}$; the lower-right block, corresponding to the neurons, is related to the Hessian of an i.i.d. network of nonlinear neurons (Eq. 52). They are coupled through the simple off-diagonal blocks, reflecting the interaction between neurons and latent variables. Finally, we obtain for the four-point function:
$\Psi_{\star}^{\phi}=\frac{1+\frac{1}{\alpha\text{PR}^{D}}|g_{\text{eff}}^{2}S_{12}^{\phi}|^{2}}{|1-g_{\text{eff}}^{2}S_{12}^{\phi}|^{2}}C_{12}^{\phi}.$ (76)
@@
To obtain $\Psi^{x}_{\star}$, we add $-\frac{1}{2}x\mathcal{J}x$ to the neuronal single-site path integral (Eq. 61) and compute
$\frac{\delta^{2}S}{\delta\mathcal{J}^{2}}=-C_{12}^{x},\>\>\>\frac{\delta^{2}S}{\delta\mathcal{J}\delta\bm{C}}=\begin{pmatrix}0\\ 0\\ S_{12}^{x}\\ -C_{12}^{x\phi}\\ \end{pmatrix}.$ (77)
Inserting these into Eq. 54 and using a formula (Eq. 149) to invert the Hessian, we obtain
$\Psi_{\star}^{x}=C^{x}_{12}+\left(1+\frac{1}{\alpha\text{PR}^{D}}\right)|U|^{2}C^{\phi}_{12}\\ +UC^{x\phi}_{12}+\text{h.c.},$ (78)
where $U={g_{\text{eff}}^{2}S^{x}_{12}}/({1-g_{\text{eff}}^{2}S^{\phi}_{12}})$ (Eq. 21b with $g=g_{\text{eff}}$).
III.6 Interpreting the formulas
The analytic formulas for the four-point function of the random-mode model (Eq. 76 and Eq. 78) reveal several insights. Taking the limit $\alpha\text{PR}^{D}\rightarrow\infty$ for fixed $g_{\text{eff}}$, which leads to elements of $\bm{J}$ being i.i.d., recovers the known formulas for i.i.d. couplings (Eq. 20 and Eq. 21). Comparing to the i.i.d. formulas, any structure beyond i.i.d. connectivity (i.e., $1/\alpha\text{PR}^{D}$ being finite) strictly decreases the dimension of activity.
Our analysis reveals two key statistical features of the connectivity that characterize how connectivity shapes activity: the effective coupling strength, $g_{\text{eff}}$, and the effective rank, $\alpha\text{PR}^{D}$. The first characterizes the typical magnitude of couplings, while the second characterizes global information about the number of significant connectivity components. As in the i.i.d. model, $g_{\text{eff}}$ influences both single-neuron activity (described by $C^{a}_{\star}(\tau)$) and collective activity (described by $\Psi^{a}_{\star}(\tau)$). In contrast, $\alpha\text{PR}^{D}$ influences collective activity exclusively. The random-mode model therefore provides independent control over single-neuron and collective aspects of activity.
To validate the theoretical prediction for $\Psi^{\phi}_{\star}(\bm{\tau})$ (Eq. 76), we simulated networks and computed the participation ratio from the empirical equal-time covariance matrix of the activations. For the component strengths $D_{a}$, we used a spectrum given by
$D_{a}\propto\exp(-\beta_{D}a/M)$ (79)
for $a=1,\ldots,M$, which yields $\text{PR}^{D}=\beta^{-1}_{D}\tanh(\beta_{D})$. The corresponding density is $P(D)=1/\beta_{D}D$ with support on $[\exp(-\beta_{D}),1]$. To attain a desired effective rank $\alpha\text{PR}^{D}<1$, we set $\alpha=M/N=1$ and solved for $\beta$. For $\alpha\text{PR}^{D}>1$, we fixed $\text{PR}^{D}=1$ and increased $\alpha$. We adjusted $g_{\text{eff}}$ by rescaling $D_{a}$ (i.e., adjusting the proportionality constant in Eq. 79).
Our simulations show excellent agreement with theoretical predictions across parameter settings (Fig. 6). Fig. 6a illustrates how the dimension of activity, $\text{PR}^{\phi}$, varies with $\alpha\text{PR}^{D}$ for different values of $g_{\text{eff}}$. Consistent with i.i.d.-coupling networks, $\text{PR}^{\phi}$ increases monotonically with $g_{\text{eff}}$ [25]. It also increases monotonically with $\alpha\text{PR}^{D}$, approaching the dimension of a network with i.i.d. couplings for large $\alpha\text{PR}^{D}$ (Fig. 6a inset). As $\alpha\text{PR}^{D}$ approaches zero, the dimension of activity also approaches zero. We examine the limits $g_{\text{eff}}\rightarrow 1^{+}$ and $\alpha\text{PR}^{D}\rightarrow 0$ in detail in Sec. IV.
Beyond the dimension of activity, which depends on $\Psi^{\phi}_{\star}(\bm{\tau})$ at $\bm{\tau}=(0,0)$, we also examined the temporal profile of this function. Fig. 6b shows a normalized version of $\Psi^{\phi}(\tau,0)$, which is related to the correlation functions of leading principal components, for various values of $g_{\text{eff}}$ and $\alpha\text{PR}^{D}$. The decay timescale of $\Psi^{\phi}_{\star}(\tau,0)$ decreases with increasing $g_{\text{eff}}$, approaching a limiting behavior at $g_{\text{eff}}\sim 10$. While $\alpha\text{PR}^{D}$ affects the overall scale of this function, and consequently the dimension of activity, it has little effect on the decay timescale, and thus little effect on the timescales of leading principal components (Fig. 6b, inset). Thus, the effective rank primarily influences the dimension of activity rather than the collective temporal structure.
IV Limiting behavior of the random-mode model
Figure 7: Analysis of dimension of activity and covariance spectrum in the low-effective-rank limit. (a) Top: Dimension of activity $\text{PR}^{\phi}$ versus effective rank $\alpha\text{PR}^{D}$ ($\alpha=1$) for various coupling strengths $g_{\text{eff}}$. Component strengths $D_{a}$ are constant for $a/N\leq\text{PR}^{D}$ and 00 otherwise. Small dots: individual simulations; large dots: means over 10 simulations; solid lines: theoretical predictions; dotted lines: small-$\text{PR}^{D}$ expansion, $K^{\phi}(g_{\text{eff}})\mathrm{PR}^{D}$. Inset: Ratio of proportionality factor $K^{\phi}(g_{\text{eff}})$ to $\text{PR}^{\phi}_{\text{i.i.d.}}$ versus $g_{\text{eff}}$. Bottom: Same analysis for pre-activations. (b) Top: Rank-ordered eigenvalue plots for activation covariance matrix $C^{\phi}_{ij}(0)$. Dotted vertical lines indicate $N\,\text{PR}^{D}$. Bottom: Same for pre-activations. All simulations use $N=6500$ neurons.
Here we examine the behavior of the random-mode model in two limiting cases: as the system approaches the transition to chaos from above ($g_{\text{eff}}\rightarrow 1^{+}$), and in the limit of low effective rank ($\alpha\text{PR}^{D}\ll 1$).
We first consider approaching the transition to chaos from above, i.e., $g_{\text{eff}}\rightarrow 1^{+}$. The relevant expansion parameter is $g_{\text{eff}}-1\ll 1$. For networks with i.i.d. couplings, Clark et al. [25] showed that, for coupling strength $g$, $\Psi^{a}_{\star,\text{i.i.d.}}(\bm{\tau})\sim 1/(g-1)$ and $\text{PR}^{a}_{\text{i.i.d.}}\sim(g-1)^{3}$ in this limit. For the random-mode model, to leading order in $1/(g_{\text{eff}}-1)$, we have
Thus, as the network approaches the transition to chaos, the effect of structure in connectivity is to reduce the dimension by a factor of $(1+1/\alpha\text{PR}^{D})^{-1}$ compared to a network with i.i.d. couplings. Taking the further limit of $\alpha\text{PR}^{D}\ll 1$ yields the simple relation
$\text{PR}^{a}=\alpha\text{PR}^{D}\text{PR}^{a}_{\text{i.i.d.}}$ (82)
The difference between $\alpha\text{PR}^{D}$ and $\text{PR}^{S}$ is a higher-order correction, so we also have
$\text{PR}^{a}=\text{PR}^{S}\text{PR}^{a}_{\text{i.i.d.}}$ (83)
That is, in the limits of both low effective rank and small effective coupling strength, the dimension of activity is equal to the dimension of the singular-value spectrum times the dimension of activity in an i.i.d. network.
We now consider the limit of low effective rank, $\alpha\text{PR}^{D}\ll 1$, for arbitrary $g_{\text{eff}}$. In this limit, the component strengths $D_{a}$ closely approximate the singular values of the coupling matrix. To leading order in $\alpha\text{PR}^{D}$,
$\text{PR}^{a}=K^{a}(g_{\text{eff}})\alpha\text{PR}^{D},$ (84)
where
@@ IV Limiting behavior of the random-mode model
[…]
We match this theory of limiting behavior to simulated networks with $\alpha=1$ and a step function for the component strengths:
$D_{a}=\begin{cases}\text{const.}&\frac{a}{N}\leq\text{PR}^{D}\\ 0&\frac{a}{N}>\text{PR}^{D}.\end{cases}$ (87)
We varied $\text{PR}^{D}$ and verified its linear relationships with $\text{PR}^{\phi}$ and $\text{PR}^{x}$ in the $\text{PR}^{D}\rightarrow 0$ limit (Fig. 7a).
Focusing on $K^{\phi}(g_{\text{eff}})$, note that its expression (Eq. 85) resembles the formula for the dimension of activity in a i.i.d. network with coupling strength $g_{\text{eff}}$, but includes an additional factor $|g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}$ in the integral (compare to Eq. 20). Numerical evaluation of $K^{\phi}(g_{\text{eff}})$ shows that it approaches $\text{PR}^{\phi}_{\text{i.i.d.}}$ as $g_{\text{eff}}\rightarrow 1^{+}$ as expected from Eq. 82. As $g_{\text{eff}}$ increases, $K^{\phi}(g_{\text{eff}})$ increases monotonically to approximately $1.53\,\text{PR}^{\phi}_{\text{i.i.d.}}$ (Fig. 7a, inset). These results show that for networks with low effective rank, $\text{PR}^{\phi}\sim\alpha\text{PR}^{D}\text{PR}^{\phi}_{\text{i.i.d.}}$ up to an order-one fudge factor that becomes exactly one as $g_{\text{eff}}\rightarrow 1^{+}$.
We also investigated the relationship between the eigenvalues $\lambda^{a}_{i}$ of $C^{a}_{ij}(0)$ and the component strengths $D_{a}$. In the low-effective-rank limit, the quantities $\text{PR}^{a}$ and $\alpha\text{PR}^{D}$ computed from these spectra exhibit a linear relationship (with proportionality factor $K^{a}(g_{\text{eff}})$), implying that the ratio of their second-squared and fourth moments are tightly linked. However, we find that the full forms of the eigenvalue and component-strength spectra differ markedly. Whereas the component strengths $D_{a}$ decay according to a step function in this example, the rank-ordered eigenvalues $\lambda^{\phi}$ and $\lambda^{x}$ decay smoothly (Fig. 7b). Thus, the dynamics transform the connectivity spectrum into an activity spectrum with a different form while preserving the linear relationship between $\text{PR}^{a}$ and $\alpha\text{PR}^{D}$.
V Cavity analysis of the random-mode model
While the path-integral approach provides a systematic framework for analyzing neural-network dynamics, the cavity method offers an intuitive, complementary perspective. This section provides an overview of the cavity analysis for the random-mode model, with the full mathematical details available in Appendix G. While these approaches yield equivalent results, they differ in their methodology: the path-integral approach focuses on fluctuations around the saddle point in a field theory, whereas the cavity approach examines perturbations induced by a subset of held-out variables.
Leveraging Eq. 56, we reformulate the network as a bipartite system of neurons and latent variables and extend the cavity calculation of Clark et al. [25] to both groups. This leads to separate, mutually referential cavity pictures for each group, mirroring the two single-site path integrals (Sec. III.5). As in the path-integral formalism, self-consistent expressions in one picture are defined using averages from the other. Unlike the path-integral approach, which derives $\Psi^{\phi}(\bm{\tau})$ using its time-by-time definition (Eq. 9), the cavity method derives this function using its neuron-by-neuron definition (Eq. 7).
The cavity method we employ, schematized in Fig. 8, has two key features that distinguish it from simpler cavity calculations. First, it uses a two-site structure: we consider the simultaneous removal of two neurons, or two latent variables, from the network, rather than just one. This allows us to study pairwise correlations between neurons. This approach was previously used in Clark et al. [25]. Second, it incorporates a bipartite structure: we treat both neurons and latent variables as dynamic objects, creating separate but mutually referential cavity pictures for each group. Such bipartite structure, without the two-site structure, has been used in cavity calculations in the Hopfield model [61, 62]. Note that these two features are distinct: the two-site aspect refers to the number of neurons removed in each cavity, while the bipartite aspect refers to the types of variables considered (neurons and latent variables).
VI Single-neuron heterogeneity
While the random-mode model introduces heterogeneity in connectivity through the component strengths $D_{a}$, individual neurons may also exhibit heterogeneous properties. For instance, cortical neurons display a broad distribution of mean firing rates [63]. To compare the effects of low-dimensional structure in connectivity with the effects of heterogeneous neuronal properties, we extend our analysis to networks with diverse single-neuron characteristics.
In particular, we generalize the neuronal nonlinearity to $\Phi_{\bm{\theta}}(x)$, which depends on a vector of parameters $\bm{\theta}$. The parameter vector for the $i$-th neuron, $\bm{\theta}_{i}$, is sampled i.i.d. from a distribution $P(\bm{\theta})$. This formulation allows us to model diverse neuronal properties such as gains or biases. When analyzing neural data, it is common to normalize neuronal activities (e.g., through $z$-scoring) to put them on the same scale. To facilitate analysis of such normalized activations, we also introduce a set of variables $\phi_{i}(t)=\phi(x_{i}(t))$ that do not depend on $\bm{\theta}_{i}$.
We incorporate the generalized nonlinearity into the path integral, as well as a source term for the normalized variables, and integrate out $\bm{L}$ and $\bm{R}$. To factorize the action over indices spanning extensive dimensions, we introduce, among others, the order parameter
$C^{\Phi}(t_{1},t_{2})=\frac{1}{N}\sum_{i}\Phi_{\bm{\theta}_{i}}(t_{1})\Phi_{\bm{\theta}_{i}}(t_{2}),$ (88)
where $\Phi_{\bm{\theta}_{i}}(t)=\Phi_{\bm{\theta}_{i}}(x_{i}(t))$. In Appendix H, we derive a single-site dynamic process $T[x](t)=\eta^{x}(t)$, where $\eta^{x}\sim\mathcal{GP}(0,\alpha r_{2}C^{\Phi}_{\star})$, with the self-consistency condition
$C^{\Phi}_{\star}(t_{1},t_{2})=\left\langle\left\langle\Phi_{\bm{\theta}}(t_{1})\Phi_{\bm{\theta}}(t_{2})\right\rangle_{\star}\right\rangle_{\bm{\theta}}.$ (89)

### v2

@@ F.3 Adding sources
Our goal is to compute
@@
where
$\tilde{\mathcal{S}}[\bm{C},U,\hat{U}]=\hat{U}U+\mathcal{S}[\bm{C};U].$ (93)
where
The fluctuations of $\hat{U}$ are given by the bottom-right element of the inverse Hessian evaluated at the saddle point.
Consider a general intensive action $\mathcal{S}[\bm{C};\mathcal{J}]$ where $\bm{C}$ is a collection of fields (e.g., $\bm{C}=\{C^{\phi},\hat{C}^{\phi}\}$), and $\mathcal{J}$ is a source. Using a Schur complement to compute this element, we obtain
$\left.\frac{\delta^{2}}{\delta\mathcal{J}^{2}}\log Z\right|_{\mathcal{J}=0}=\left[-\frac{\delta^{2}S}{\delta\mathcal{J}^{2}}+\left(\frac{\delta^{2}S}{\delta\mathcal{J}\delta\bm{C}}\right)^{T}\bm{H}^{-1}\frac{\delta^{2}S}{\delta\mathcal{J}\delta\bm{C}}\right]_{\mathcal{J}=0},$ (96)
with all quantities evaluated at the saddle point. By applying this formula to the action of the i.i.d. model with the source term $\frac{1}{2}x\mathcal{J}x$ added to the single-site path integral (Eq. 76), we recover the expression for $\Psi^{x}_{\star}(\bm{\tau})$ from the two-site cavity method (Clark et al. [30]; Eq. 73). For actions of the form $\mathcal{S}[\bm{C};\mathcal{J}]=-\mathcal{J}C_{a}+\text{$\mathcal{J}$-independent terms}$, this formula reduces to $({\delta^{2}}/{\delta\mathcal{J}^{2}})\log Z|_{\mathcal{J}=0}=[\bm{H}^{-1}]_{aa},$ as expected.
$\frac{\delta^{2}S}{\delta\mathcal{J}^{2}}=-C_{12}^{x},\quad\frac{\delta^{2}S}{\delta\mathcal{J}\delta\bm{C}}=\begin{pmatrix}g^{2}\left(S_{12}^{x}\right)^{*}\\ -C_{12}^{x\phi}\end{pmatrix}$, (98)
F.4 Other kinds of path integrals
Different path-integral formalisms have been applied to neural network models depending on the underlying dynamics. The Doi-Peliti formalism [93] derives path integrals from master equations for systems with non-negative integer degrees of freedom (such as spiking neural networks) that evolve according to master equations, using an operator formalism as an intermediate step. In contrast, we use the MSRDJ formalism for continuous-time stochastic differential equations, which naturally describes our rate-based network dynamics. For a comprehensive discussion of different path-integral approaches in neural network theory, see [60].
Appendix G Temporal limits and inverses
We start by examining the properties of $(t_{1},t_{2},t_{3},t_{4}). UU$ vanishes if one of the four time points is far (measured on the timescale of $C^{\phi}_{\star}(\tau)$) from the other three. For $t_{1}$ or $t_{2}$, this is a consequence of the perturbation associated with $\hat{x}(t_{1})$ or $\hat{x}(t_{2})$ being far away; for $t_{3}$ or $t_{4}$, this is a consequence of $\langle\phi(t_{i})\rangle_{\star}=0$ separating from the expectation. $U$ also vanishes if there is more than one time point which is far from the remaining ones by the same arguments. Thus, the time points need to be close in pairs. However, if $t_{1},t_{2}$ are close and $t_{3},t_{4}$ are close, but the pairs are distant, $U$ still vanishes since the perturbations associated with $\hat{x}(t_{i})$ are both far. This leaves three possibilities for a non-vanishing $U$: $t_{1},t_{3}$ close and $t_{2},t_{4}$ close, but the pairs are distant; $t_{1},t_{4}$ close and $t_{2},t_{3}$ close, but the pairs are distant; or all time points are close.
This justifies taking the limits Eq. 34 for $K$.
Appendix H Partially symmetric disorder
$\left\langle J_{ij}J_{k\ell}\right\rangle=g^{2}\delta_{ik}\delta_{j\ell}+g^{2}\rho\delta_{i\ell}\delta_{jk}.$ (105)
$Z[\bm{X},\bm{Y}]=\int\mathcal{D}\bm{x}\int\mathcal{D}\hat{\bm{x}}\exp\bigg{(}i\sum_{i=1}^{N}\hat{x}_{i}T[x_{i}]-i\sqrt{{g^{2}-|\sigma|}}\sum_{ij}X_{ij}\hat{x}_{i}\phi_{j}-i\sqrt{\frac{|\sigma|}{2}}\sum_{ij}Y_{ij}(\hat{x}_{i}\phi_{j}+\text{sgn}(\sigma)\phi_{i}\hat{x}_{j})\bigg{)}.$ (107)
$C^{\phi}(t_{1},t_{2})$ $=\frac{1}{N}\sum_{i=1}^{N}\phi_{i}(t_{1})\phi_{i}(t_{2}),$ $\hskip-72.26999pt\text{conjugate: }\hat{C}^{\phi}(t_{1},t_{2});$ (108)
$S^{\phi}(t_{1},t_{2})$ $=\frac{1}{N}\sum_{i=1}^{N}\frac{\delta\phi_{i}(t_{1})}{\delta I_{i}(t_{2})},$ $\hskip-72.26999pt\text{conjugate: }\hat{S}^{\phi}(t_{1},t_{2}).$ (109)
We obtain, for the disorder-averaged path integral,
H.1 Two-point functions
$\hat{C}^{\phi}_{\star}(t_{1},t_{2})$ $=0,$ (117)
$\eta^{x}$ $\sim\mathcal{GP}(0,g^{2}C^{\phi}_{\star}),$ (121)
H.2 Four-point functions
$\bm{H}=\begin{pmatrix}0&-1+g^{2}S^{\phi}_{12}&0&0\\ -1+g^{2}(S^{\phi}_{12})^{*}&-C^{\phi}_{12}&\sigma(S^{\phi}_{1})^{*}C^{\phi}_{2}&\sigma(S^{\phi}_{2})^{*}C^{\phi}_{1}\\ 0&\sigma S^{\phi}_{1}C^{\phi}_{2}&0&\sigma-\sigma^{2}S^{\phi}_{1}(S^{\phi}_{2})^{*}\\ 0&\sigma S^{\phi}_{2}C^{\phi}_{1}&\sigma-\sigma^{2}(S^{\phi}_{1})^{*}S^{\phi}_{2}&0\end{pmatrix}$. (122)
Appendix I Path-integral calculation of two- and four-point functions for the random-mode model
Having demonstrated the path-integral approach to computing two- and four-point functions for i.i.d. couplings, we now apply this formalism to the random-mode model. In terms of the mode matrices $\bm{L}$ and $\bm{R}$, the path integral is
$Z[\bm{L},\bm{R}]=\int\mathcal{D}\bm{x}\int\mathcal{D}\hat{\bm{x}}\exp\left(i\sum_{i=1}^{N}\hat{x}_{i}T[x_{i}]-i\sum_{i,j=1}^{N}\left[\bm{L}\bm{D}\bm{R}\right]_{ij}\hat{x}_{i}\phi_{j}\right).$ (124)
@@ Appendix I Path-integral calculation of two- and four-point functions for the random-mode model
$Z[\bm{L},\bm{R}]=\int\mathcal{D}\bm{x}\int\mathcal{D}\hat{\bm{x}}\int\mathcal{D}\bm{z}\int\mathcal{D}\hat{\bm{z}}\exp\Big{(}i\sum_{i=1}^{N}\hat{x}_{i}T[x_{i}]+i\sum_{a=1}^{M}z_{a}\hat{z}_{a}-i\sum_{i,a}L_{ia}\hat{x}_{i}z_{a}-i\sum_{i,a}D_{a}R_{ia}\phi_{i}\hat{z}_{a}\Big{)}.$ (126)
@@
[…]
@@
[…]
@@
I.1 Two-point functions
Following the steps outlined at the end of Sec. F.2, we first compute the DMFT by calculating the saddle point. The derivatives of the action are
@@ I.1 Two-point functions
Solving for $Q_{\star}(t_{1},t_{2})$ by combining Eq. 138, Eq. 140, and Eq. 142, we obtain $Q_{\star}(t_{1},t_{2})=\alpha r_{2}C^{\phi}_{\star}(t_{1},t_{2}),$ with $r_{n}$ defined by Eq. 10. Consolidating these results yields Eq. 36c.
I.2 Four-point functions
Figure 10: Analysis of dimension of activity and covariance spectrum in the low-effective-rank limit. (a) Top: Dimension of activity $\text{PR}^{\phi}$ versus effective rank $\alpha\text{PR}^{D}$ ($\alpha=1$) for various coupling strengths $g_{\text{eff}}$. Component strengths $D_{a}$ are constant for $a/N\leq\text{PR}^{D}$ and 0 otherwise. Small dots: individual simulations; large dots: means over 10 simulations; solid lines: theoretical predictions; dotted lines: small-$\text{PR}^{D}$ expansion, $K^{\phi}(g_{\text{eff}})\mathrm{PR}^{D}$. Inset: Ratio of proportionality factor $K^{\phi}(g_{\text{eff}})$ to $\text{PR}^{\phi}_{\text{i.i.d.}}$ versus $g_{\text{eff}}$. Bottom: Same analysis for pre-activations. (b) Top: Rank-ordered eigenvalue plots for activation covariance matrix $C^{\phi}_{ij}(0)$. Dotted vertical lines indicate $N\,\text{PR}^{D}$. Bottom: Same for pre-activations. All simulations use $N=6500$ neurons.
@@ I.2 Four-point functions
$\bm{H}=\left(\begin{array}[]{cc|cc}0&\alpha r_{2}&0&-1\\ \alpha r_{2}&-\alpha r_{4}C_{12}^{\phi}&-1&0\\ \hline\cr 0&-1&0&S_{12}^{\phi}\\ -1&0&(S_{12}^{\phi})^{*}&-C_{12}^{\phi}\end{array}\right).$ (143)
@@
The upper-left block, corresponding to the latent variables, is related to the Hessian of an i.i.d. network with linear dynamics and a heterogeneous distribution of single-neuron scale factors $D_{a}$; the lower-right block, corresponding to the neurons, is related to the Hessian of an i.i.d. network of nonlinear neurons (Eq. 88). They are coupled through the simple off-diagonal blocks, reflecting the interaction between neurons and latent variables.
@@
This yields the expression for $\Psi^{\phi}_{\star}$ (Eq. 38). To obtain $\Psi^{x}_{\star}$, we add $\frac{1}{2}x\mathcal{J}x$ to the neuronal single-site path integral (Eq. 130) and compute
$\frac{\delta^{2}S}{\delta\mathcal{J}^{2}}=-C_{12}^{x},\>\>\>\frac{\delta^{2}S}{\delta\mathcal{J}\delta\bm{C}}=\begin{pmatrix}0\\ 0\\ (S_{12}^{x})^{*}\\ -C_{12}^{x\phi}\\ \end{pmatrix}$. (144)
Inserting these into Eq. 96 we obtain
$\Psi_{\star}^{x}=C^{x}_{12}+\left(1+\frac{1}{\alpha\text{PR}^{D}}\right)|U|^{2}C^{\phi}_{12}+UC^{x\phi}_{12}+\text{h.c.},$ (145)
where $U={g_{\text{eff}}^{2}S^{x}_{12}}/({1-g_{\text{eff}}^{2}S^{\phi}_{12}})$ (Eq. 73b with $g=g_{\text{eff}}$). Here, we have used the following formula to invert the Hessian:
$\begin{pmatrix}0&a&0&-1\\ a^{*}&b&-1&0\\ 0&-1&0&c\\ -1&0&c^{*}&d\end{pmatrix}^{-1}=\begin{pmatrix}-\frac{b+|c|^{2}d}{|1-ac|^{2}}&-\frac{c^{*}}{1-(ac)^{*}}&-\frac{ad+bc^{*}}{|1-ac|^{2}}&-\frac{1}{1-(ac)^{*}}\\ -\frac{c}{1-ac}&0&-\frac{1}{1-ac}&0\\ -\frac{a^{*}d+bc}{|1-ac|^{2}}&-\frac{1}{1-(ac)^{*}}&-\frac{|a|^{2}d+b}{|1-ac|^{2}}&-\frac{a^{*}}{1-(ac)^{*}}\\ -\frac{1}{1-ac}&0&-\frac{a}{1-ac}&0\end{pmatrix}$. (146)
Appendix J Path-integral calculation of two- and four-point functions for the random-mode model with single-unit heterogeneity
We incorporate the generalized nonlinearity $\Phi_{\bm{\theta}_{i}}(x)$ into the path integral, as well as a source term for the normalized variables, and integrate out $\bm{L}$ and $\bm{R}$. To factorize the action over indices spanning extensive dimensions, we introduce the order parameters
$C^{\Phi}(t_{1},t_{2})$ $=\frac{1}{N}\sum_{i=1}^{N}\Phi_{\bm{\theta}_{i}}(t_{1})\Phi_{\bm{\theta}_{i}}(t_{2}),$ $\hskip-72.26999pt\text{conjugate: }\hat{Q}(t_{1},t_{2});$ (147)
$Q(t_{1},t_{2})$ $=\frac{1}{N}\sum_{a=1}^{M}z_{a}(t_{1})z_{a}(t_{2}),$ $\hskip-72.26999pt\text{conjugate: }\hat{C}^{\phi}(t_{1},t_{2}),$ (148)
where $\Phi_{\bm{\theta}_{i}}(t)=\Phi_{\bm{\theta}_{i}}(x_{i}(t))$. The resulting action and single-site path integrals for this system are
@@
J.1 Two-point functions
$\hat{C}^{\Phi}_{\star}(t_{1},t_{2})$ $=\hat{Q}_{\star}(t_{1},t_{2})=0,$ (152)
$C^{\Phi}_{\star}(t_{1},t_{2})$ $=\left\langle\left\langle\Phi_{\bm{\theta}}(t_{1})\Phi_{\bm{\theta}}(t_{2})\right\rangle_{\star}\right\rangle_{\bm{\theta}},$ (153)
where $\left\langle\cdots\right\rangle_{\star}$ and $\left\langle\cdots\right\rangle_{\star D}$ are averages with respect to $W^{x}_{\bm{\theta}}[Q_{\star},0;0]$ and $W^{z}_{D}[C^{\Phi}_{\star},0]$, respectively (note that $W^{x}_{\bm{\theta}}[Q_{\star},0;0]$ does not depend on $\bm{\theta}$). Using $\left\langle z(t_{1})z(t_{2})\right\rangle_{\star D}=D^{2}C^{\Phi}_{\star}(t_{1},t_{2})$ and, consequently, $Q_{\star}(t_{1},t_{2})=\alpha r_{2}C^{\Phi}_{\star}(t_{1},t_{2})$, we obtain the self-consistent single-site dynamic process
$T[x](t)$ $=\eta^{x}(t),$ (155a)
$\eta^{x}$ $\sim\mathcal{GP}(0,\alpha r_{2}C^{\Phi}_{\star}),$ (155b)
$C^{\Phi}_{\star}(t_{1},t_{2})$ $=\left\langle\left\langle\Phi_{\bm{\theta}}(t_{1})\Phi_{\bm{\theta}}(t_{2})\right\rangle_{\star}\right\rangle_{\bm{\theta}}.$ (155c)